Production function of a firm is given as X=0.5 L ^0.5 K^0.5 price of labor and capital are 5$ and 10$ respectively the firm has a constant cost out lay of $600 find the combination of labor and capital that maximize the farm output and the maximum outputs?

1
Expert's answer
2022-02-06T14:23:02-0500

"MPL=\frac{K^{0.5}}{4L^{0.5}}"


"MPK=\frac{L^{0.5}}{4K^{0.5}}"


"MRTS = \frac{MPL} {MPK}"


"MRTS = \frac{ \frac{K^{0.5}}{4L^{0.5}}}{\frac{L^{0.5}}{4K^{0.5}}}"


"MRTS = \frac{K}{L}"


"\frac{K}{L}= \frac{10}{5}"


10L=5K


2L=K


"600=0.5L^{0.5} K^{0.5}"


"600=0.5L^{0.5} (2L)^{0.5}"


"1200=L^{0.5}L^{0.5}2^{0.5}"


"1200=(2)^{0.5}L"


L=848.53


K=1697.06